1) In an economy where people spend 75% of their incomes, the government increases taxes by $6000. Calculate the change in real GDP that results.

(round your final answer to two decimal points. Use a positive number to denote an increase and a negative number (with negative sign) to denote a decrease.)

2) In an economy where people spend 75% of their incomes, the government lowers its spending by $6000. Calculate the change in real GDP that results.

(round your final answer to two decimal points. Use a positive number to denote an increase and a negative number (with negative sign) to denote a decrease.)
Answer #1

1) MpC=75%=0.75

then tax multiplier=-Mpc/1-Mpc=0.75/1-0.75= -3

Increase in taxes will decrease real GDP by 18000

Change in GDP=-18000

2)Govt expenditure multiplier=1/1-Mpc=1/1-0.75=4

Decrease in G will decrease real GDP by 24000